OpenStudy (ashking1):

Find the real numbers x and y that make the equation true. -4 + yi = x + 3i

OpenStudy (freckles):

if x=-4 then x is ? if yi=3i then y is?

OpenStudy (freckles):

for example if a+bi=c+di then we have a=c which means a is c and we have bi=di which means b=d or b is d

OpenStudy (ashking1):

so they swap positions its actually -4 +3i = -4+ 3i

OpenStudy (freckles):

can you say what x and y are? like what is in place of x and what is in place of y?

OpenStudy (freckles):